OpenStudy (anonymous):

how do i solve this 16 < 2x + 4 < 4 ?

OpenStudy (johnweldon1993):

You want to isolate the 2x in the middle first...so subtract 4 from all sides of the equation \[\large 16 - 4 < 2x + 4 -4 < 4 - 4\] this now becomes \[\large 12 < 2x < 0\] Now you want to solve for 'x'...not 2x....so divide everything by 2... \[\large \frac{12}{2} < \frac{2x}{2} < \frac{0}{2}\] And this becomes... \[\large ? < x < \space ?\]